OpenStudy (anonymous):

2x^2+x-6=0 I got to the square rooting then lost it. My options are: (-2,3/2) (-3/2,2) (0)

OpenStudy (souvik):

you have to factor it

OpenStudy (anonymous):

What do you mean factor? Isn't just like the last problem?

OpenStudy (souvik):

no its not like the previous one...!

OpenStudy (anonymous):

Damn..

OpenStudy (souvik):

let me sure...is it like..\[\huge\color\red{2x ^{2}+x-6=0}\]

OpenStudy (anonymous):

Yes

OpenStudy (souvik):

do u know how to factor?

OpenStudy (anonymous):

Not in this equation.

OpenStudy (souvik):

so your teacher does no teach u this....!

OpenStudy (anonymous):

Nope! Because this is a online class. I don't have a teacher.

OpenStudy (uri):

So you don't have online teachers? :3

OpenStudy (anonymous):

No.

OpenStudy (uri):

Oh my my :o anyway @souvik will help you :)

OpenStudy (anonymous):

Okay(: Thank you(:

OpenStudy (souvik):

\[2x^2+x-6=0\] or 2x^2+4x-3x-6=0 or,2x(x+2)-3(x+2)=0 or, (x+2)(2x-3)=0 so (x+2)=0 or x=-2 and (2x-3)=0 or x=3/2

OpenStudy (anonymous):

Thank you(:

OpenStudy (souvik):

welcome..:)